OpenStudy (anonymous):

Simplify the trigonometric expression: sin 0 cos 0

OpenStudy (anonymous):

\[\sin \theta \cos \theta \]

OpenStudy (agent0smith):

\[\large 2\sin \theta \cos \theta = \sin 2 \theta\]so \[\large \sin \theta \cos \theta = \frac{ 1 }{ 2 }\sin 2 \theta\]

OpenStudy (anonymous):

but ...that none of my answers :P

OpenStudy (anonymous):

a. \[\cot \theta\]

OpenStudy (anonymous):

b. 1

OpenStudy (anonymous):

c. \[\sin \theta \]

OpenStudy (anonymous):

d. \[\tan \theta \]

OpenStudy (agent0smith):

None of those are valid options for sinθcosθ

OpenStudy (anonymous):

ok then haha cx